OpenStudy (anonymous):

Suppose that $80,000 is invested at 6% interest. Find the amount of money in the account after 7 years if the interest is compound annually. If interest is compounded annually, what is the amount of money after t=7 years? $____

OpenStudy (anonymous):

fv = pv(1+i)^n fv = 80000(1+0.06)^7 fv = 120290.42

OpenStudy (anonymous):

Ok one more :)

OpenStudy (anonymous):

kk